# Aerial Vision-Language Navigation with a Unified Framework for Spatial, Temporal and Embodied Reasoning

### v1

@@ III-C Framework Overview
Long-horizon trajectories may yield prohibitively long visual token sequences. To mitigate this, we apply a Spatial Token Compression (STC) module that aggregates non-overlapping $g\times g$ neighborhoods. A lightweight MLP projector aligns visual features to the LLM embedding space.
$\tilde{E}^{\text{vis}}_{t_{i}}$ $=\mathrm{STC}(E^{\text{vis}}_{t_{i}})\in\mathbb{R}^{(N/g^{2})\times(Cg^{2})},$ (2)
$Z^{\text{vis}}_{t_{i}}$ $=\mathrm{Proj}_{\text{mlp}}(\tilde{E}^{\text{vis}}_{t_{i}})\in\mathbb{R}^{(N/g^{2})\times D}.$
@@
$S_{t}=[Z^{\text{vis}}_{0},\,\ldots,\,Z^{\text{vis}}_{t_{K}},\,Z^{\text{text}}],$ (3)
@@
Action Parsing and Execution. Once the model outputs a textual action command (e.g., “The next action is move forward 15 units”), we apply a regular-expression parser to extract the action type and its numeric argument. The parsed action is then decomposed into low-level motion primitives (e.g., multiple fixed-step forward movements). These primitives are executed in an open-loop manner until all of them are completed, after which the agent receives the next visual observation from the environment, updates its trajectory history, and predicts the subsequent action.
@@ III-D Training Objective
Our model is trained in a multi-task setting that jointly supervises aerial navigation, spatial perception, and trajectory reasoning through autoregressive language modeling. Since navigation trajectories exhibit a highly imbalanced distribution of actions, we apply a frequency-based label reweighting strategy. Each action $a\in\mathcal{A}$ is assigned a normalized inverse-frequency weight,
@@
The overall multi-task objective integrates supervision from all three tasks. Let $\mathcal{B}$ denote the mixed training set. The loss is defined as
@@ III-E Implementation Details
Dataset Curation. To support the multi-task training objectives, we curate a mixed dataset that provides complementary supervision for action prediction, spatial grounding, and temporal reasoning, shown in Table I. Our curation is guided by two principles: maintaining a coherent input-output structure across tasks and constructing a balanced mixture that captures the diversity of aerial navigation scenarios. For aerial navigation task, we derive step-wise supervision from AerialVLN simulation trajectories by applying action merging and keyframe selection, producing compact training tuples $(\tau_{t},I,a_{t})$ . For spatial perception, we incorporate VQA-style supervision from the drone-centric subset of Open3D-VQA  [zhang2025open3dvqa], complemented by a spatial-relation-focused subset of GQA [hudson2019gqa] filtered for geometric predicates. To provide temporal abstraction signals, trajectory reasoning annotations are obtained by segmenting AerialVLN trajectories into ordered sub-trajectories and generating progress summaries aligned with instruction semantics. This curated mixture forms a coherent and comprehensive data blend for supervised fine-tuning of the proposed aerial VLN model.
Model Training. We fine-tune NVILA-lite-2B and NVILA-lite-8B on our aerial navigation task, initializing both models from their pretrained weights. Training is performed on 4 NVIDIA A100 GPUs using DeepSpeed ZeRO-3 optimization, with a learning rate of $2\times 10^{-5}$. We employ a cosine decay schedule with a warm-up ratio of 0.03. We adopt a global batch size of 128, achieved through a per-GPU batch size of 8 and 4 gradient accumulation steps. All models are trained for 1 epoch.
Method Observation Validation Seen Validation Unseen
[…]
Grid-based VS [zhao2025aerialviewselection] $\checkmark$ $\checkmark$ 70.3 20.8 33.4 10.2 121.3 7.4 16.1 2.5
CityNavAgent [zhang2025citynavagent] $\checkmark$ $\checkmark$ $\checkmark$ 80.8 13.9 30.2 5.1 60.2 11.7 35.2 5.0
Random Sampling 109.6 0.0 0.0 0.0 149.7 0.0 0.0 0.0
Action Sampling 213.8 0.9 5.7 0.3 237.6 0.2 1.1 0.1
LingUNet [misra2018mapping] $\checkmark$ $\checkmark$ 383.8 0.6 6.9 0.2 368.4 0.4 3.6 0.9
Seq2Seq [krantz2020cma] $\checkmark$ $\checkmark$ 146.0 4.8 19.8 1.6 218.9 2.3 11.7 0.7
CMA [krantz2020cma] $\checkmark$ $\checkmark$ 121.0 3.0 23.2 0.6 172.1 3.2 16.0 1.1
LAG [liu2023aerialvln] $\checkmark$ $\checkmark$ 90.2 7.2 15.7 2.4 127.9 5.1 10.5 1.4
SMTR [gao2024aerialstmr] $\checkmark$ $\checkmark$ 96.3 12.6 31.6 2.7 119.5 10.8 23.0 1.9
MapGPT [chen2024mapgpt] $\checkmark$ 124.9 2.1 4.7 0.8 107.0 0.0 0.0 0.0
Navid [gao2025openfly] $\checkmark$ 105.1 6.8 15.5 1.1 106.9 6.1 12.3 0.7
Openfly [zhang2024navid] $\checkmark$ 127.2 8.1 21.8 1.6 113.8 7.6 18.2 1.5
Ours $\checkmark$ 79.6 11.4 37.7 6.3 95.8 8.1 28.9 2.2
@@ IV-A Experimental Setup
Simulated environments. We conduct experiments on the AerialVLN-S benchmark [liu2023aerialvln], which provides a realistic aerial navigation setting rendered in AirSim and Unreal Engine 4. The dataset covers 17 compact outdoor scenes representing diverse urban environments such as residential areas, industrial zones, parks, and villages. For evaluation, the dataset allocates 12 environments to the training and validation-seen splits, while the validation-unseen split is drawn from 5 distinct scenes. Flight trajectories are collected by expert pilots, who follow naturalistic aerial routes to ensure realistic motion patterns and diverse path configurations. Overall, AerialVLN-S provides 10,113 language instructions for training, with 333 and 531 instructions assigned to the validation-seen and validation-unseen splits, respectively.
Evaluation Metrics. We adopt four standard metrics to evaluate aerial vision-and-language navigation[liu2023aerialvln]:
@@
Baselines. We compare our approach against previous diverse baselines. These baselines fall into three categories [zhang2025citynavagent]:
@@
  • Zero-shot LLM-based methods. MapGPT [chen2024mapgpt] introduces a language-based online map representation to equip GPT-4v [achiam2023gpt] with global spatial understanding, enabling adaptive multi-step planning for indoor VLN in a zero-shot setting. STMR[gao2024aerialstmr] enhances spatial reasoning by introducing a Semantic–Topo–Metric Representation that allows LLMs to infer navigation decisions directly from structured scene information. CityNavAgent[zhang2025citynavagent] performs hierarchical semantic planning to predict waypoints in a zero-shot manner, supported by a global memory module that records past trajectories for improved long-horizon navigation.
@@
  • Learning-based methods. Seq2Seq[krantz2020cma] predicts actions using a recurrent policy that encodes visual observations and language instructions into a unified hidden representation. CMA[krantz2020cma] introduces cross-modal attention to better fuse RGB, depth, and textual cues for action prediction. LAG[liu2023aerialvln] adds a look-ahead guidance module. NaVid[zhang2024navid] is a video-based vision–language model that has demonstrated strong performance in indoor VLN. OpenFly[gao2025openfly] adopts a keyframe-aware VLN architecture and uses adaptive token sampling over historical observations.
IV-B Results on AerialVLN-S Benchmark
IV-B1 Quantitative Results
To ensure a fair comparison, we categorize methods by their sensor configurations, distinguishing RGB-only agents from those that rely on panoramic inputs, depth sensing, or odometry. Within the RGB-only group, our method achieves strong performance on all metrics and attains the best overall results, with comparable to STMR [gao2024aerialstmr], which required external depth information. The proposed method is also competitive with panoramic baselines, especially in SDTW. This is particularly meaningful on AerialVLN setting, where many trajectories contain turn-backs and repeated segments, making SDTW a more informative measure of instruction following than SR alone. Panoramic methods typically require multiple cameras or repeated in-place rotations to build 360° observations, which increases time and energy costs for UAVs, whereas our model operates from a single-view RGB stream and thus better matches realistic deployment constraints. MapGPT [chen2024mapgpt] and Navid [zhang2024navid], which are strong indoor VLN baseline, perform notably worse in outdoor aerial setting, underscoring the additional challenges posed by large-scale, sparsely structured environments. Compared with the most related OpenFly [gao2025openfly] framework, our approach further benefits from long-horizon history modeling and auxiliary tasks, leading to more stable long-range behavior. Overall, these results establish a new state of the art for monocular RGB-only aerial VLN and validate the effectiveness of our unified framework for spatial, temporal, and embodied reasoning. Fig. 5 compares the predicted 3D flight paths with the reference routes. The predicted trajectories preserve the overall geometric structure of the routes, including the major turning segments, long-range direction changes, and the characteristic height variations implied by the instruction. Although small local deviations occasionally appear, the agent often recovers from these offsets and continues to produce geometrically coherent paths over extended horizons. The trajectories also converge toward the intended final region, reflecting stable motion patterns derived solely from egocentric RGB observations in large-scale outdoor environments.
@@ IV-B1 Quantitative Results
IV-B2 Qualitative Analysis
@@
Figure 7: Qualitative results on auxiliary VQA tasks. Top: examples from the Open3D-VQA dataset [zhang2025open3dvqa], where the model answers 3D spatial queries including relative relations and geometric orientation. Bottom: examples from GQA [hudson2019gqa], covering object grounding and relational reasoning. These examples illustrate the model’s of spatial understanding and positional reasoning, which is crucial for language-guided UAV navigation.
@@
IV-C Ablation Study
IV-C1 Effect of Auxiliary Tasks
We conduct ablation experiments to evaluate the effect of the two auxiliary tasks: Spatial Prediction (SP) and Trajectory Reasoning (TR). Without auxiliary supervision, the policy shows limited spatial grounding and produces the lowest SR and SDTW on both splits. SP leads to consistent gains, confirming that geometric cues improve local action grounding. Incorporating TR further boosts performance, with the largest improvements on unseen environments, where long-horizon cues are critical for handling novel layouts. These results demonstrate that the two objectives are complementary: SP enhances local spatial understanding, while TR strengthens temporal reasoning. Their combination yields a more reliable navigation policy and supports the design of our multi-task framework, reinforcing our goal of unifying spatial, temporal, and embodied reasoning within a single model.
TABLE III: Ablation study on auxiliary tasks: Spatial Prediction (SP) and Trajectory Reasoning (TR). Both tasks improve success rate (SR) and trajectory quality (SDTW) on both seen and unseen environments.

### v2

@@ III-C Framework Overview
Figure 4: Schematic diagram of the STC module with grid size $g$ = 2.
@@
Long-horizon trajectories may yield prohibitively long visual token sequences. To reduce the length of token sequences while preserving local spatial context, we introduce a Spatial Token Compression (STC) module, as illustrated in Fig. 4. Given patch-level visual tokens $E^{\text{vis}}_{t_{i}}$, STC first reshapes the token sequence into a 2D feature map. The feature map is then partitioned into non-overlapping $g\times g$ grids, where $g$ denotes the grid size. When the spatial resolution is not divisible by $g$, zero-padding is applied along the corresponding dimensions before grid partitioning. Within each grid, the $g\times g$ tokens are concatenated along the channel dimension and flattened to form a compact visual representation. This yields a compressed token sequence:
$\color[rgb]{0,0,0}\tilde{E}^{\text{vis}}_{t_{i}}=\mathrm{STC}(E^{\text{vis}}_{t_{i}})\in\mathbb{R}^{(N/g^{2})\times(Cg^{2})}.$ (2)
A lightweight MLP projector aligns visual features to the LLM embedding space:
$Z^{\text{vis}}_{t_{i}}=\mathrm{Proj}_{\text{mlp}}(\tilde{E}^{\text{vis}}_{t_{i}})\in\mathbb{R}^{(N/g^{2})\times D}.$ (3)
@@
$\color[rgb]{0,0,0}S_{t}=[Z^{\text{vis}}_{1},\,\ldots,\,Z^{\text{vis}}_{t_{K}},\,Z^{\text{text}}],$ (4)
@@
Action Parsing and Execution. During inference, the model operates with the embodied navigation-specific prompt as illustrated in Fig. 5, and directly generates textual navigation actions. Once the model outputs a textual action command (e.g., “The next action is move forward 15 units”), we apply a regular-expression parser to extract the action type and its numeric argument. The parsed action is then decomposed into low-level motion primitives (e.g., multiple fixed-step forward movements). These primitives are executed in an open-loop manner until all of them are completed, after which the agent receives the next visual observation from the environment, updates its trajectory history, and predicts the subsequent action. We quantitatively evaluated the reliability of this parsing module across simulated benchmarks, where it achieved a 100% success rate in extracting valid actions. This design follows a robust and established paradigm in vision-language navigation research, with prior works such as NaVid [44] and NaVid-4D [24] reporting equally reliable performance using the same approach in both simulation and real-world deployment.
@@ III-D Training Objective
Our model is trained in a multi-task setting that jointly supervises aerial navigation, spatial perception, and trajectory reasoning through autoregressive language modeling. Since navigation trajectories exhibit a highly imbalanced distribution of actions, we apply a frequency-based label reweighting strategy. Each action $a\in\mathcal{A}$ is assigned a normalized inverse-frequency weight.
@@
The overall multi-task objective integrates supervision from all three tasks. While $\mathcal{B}$ denotes the mixed training set, the loss is defined as
@@ III-E Implementation Details
Dataset Curation. To support the multi-task training objectives, we curate a mixed dataset that provides complementary supervision for action prediction, spatial grounding, and temporal reasoning, as shown in Table I. Our curation is guided by two principles: maintaining a coherent input-output structure across tasks and constructing a balanced mixture that captures the diversity of aerial navigation scenarios. For aerial navigation task, we derive stepwise supervision from AerialVLN simulation trajectories by applying action merging and keyframe selection, producing compact training tuples $(\tau_{t},I,a_{t})$ . For spatial perception, we incorporate VQA-style supervision from the drone-centric subset of Open3D-VQA  [46], complemented by a spatial-relation-focused subset of GQA [16] filtered for geometric predicates. To provide temporal abstraction signals, trajectory reasoning supervision is constructed by manually segmenting training trajectories into ordered sub-trajectories (41̃6 keyframes after action merging). The segmentation follows two principles: each sub-trajectory must correspond to a semantically coherent sub-instruction, and the sub-instruction must clearly describe its terminal state or landmark. Sub-instructions are directly extracted from the original instruction without paraphrasing or language model generation to preserves the original linguistic distribution. After filtering severely ambiguous instructions, approximately $3k$ trajectories are annotated, yielding around $10k$ trajectory reasoning training instances through compositional augmentation. This curated mixture forms a coherent and comprehensive data blend for supervised fine-tuning of the proposed aerial VLN model.
Model Training. We fine-tune NVILA-lite-2B and NVILA-lite-8B [26] on our aerial navigation task, initializing both models from their pretrained weights. Following NVILA, the vision encoder is implemented using SigLIP [41], the projector is a lightweight two-layer MLP that aligns visual features with the language embedding space, and the language model is based on Qwen2 [40]. Specifically, we adopt a full fine-tuning strategy, where all model components, including the vision encoder, MLP projector, and large language model, are jointly optimized, with all parameters being trainable during training. We perform supervised fine-tuning (SFT) for our tasks and the training objective is defined by the loss function presented in Eq. 7. Training is performed on 4 NVIDIA A100 GPUs using DeepSpeed ZeRO-3 optimization, with a learning rate of $2\times 10^{-5}$. We employ a cosine decay schedule with a warm-up ratio of 0.03. We adopt a global batch size of 128, achieved through a per-GPU batch size of 8 and 4 gradient accumulation steps. All models are trained for 1 epoch. For action merging, consecutive identical actions are combined, with the maximum merge length limited to 3 steps. For keyframe sampling, we uniformly sample $K=8$ frames from the visual trajectory, while ensuring that the first and the current frames are always preserved. We set the auxiliary task weights to $\lambda_{\text{sp}}=1.0$ for spatial perception and $\lambda_{\text{tr}}=0.5$ for trajectory reasoning.
@@ IV-A Experimental Setup
Simulated environments. We conduct experiments across multiple aerial VLN benchmarks for comprehensive evaluation. AerialVLN-S benchmark [25] provides a realistic aerial navigation setting rendered in AirSim and Unreal Engine 4. The dataset covers 17 compact outdoor scenes representing diverse urban environments such as residential areas, industrial zones, parks, and villages. For evaluation, the dataset allocates 12 environments to the training and validation-seen splits, while the validation-unseen split is drawn from 5 distinct scenes. Flight trajectories are collected by expert pilots, who follow naturalistic aerial routes to ensure realistic motion patterns and diverse path configurations. Overall, AerialVLN-S provides 10,113 language instructions for training, with 333 and 531 instructions assigned to the validation-seen and validation-unseen splits, respectively. In addition to AerialVLN-S, we evaluate our method on the OpenFly benchmark [12], constructed via automatic trajectory generation and vision-language-model-based instruction synthesis. OpenFly features diverse urban environments with varied scene layouts, flight heights, and trajectory lengths. Given the large scale of OpenFly, which contains over 100k navigation trajectories, we conduct experiments on a selected subset for practical evaluation. Specifically, we use trajectories collected from six high-quality UE4 scenes, forming a subset referred to as OpenFly-S. Training and evaluation are conducted on subsets comprising 61,004 and 1,210 trajectories, respectively.
Evaluation Metrics. We adopt several standard metrics to evaluate aerial vision-and-language navigation[25]:
@@
[…]
Baselines. We compare our approach against previous diverse baselines. These baselines fall into three categories [45]:
@@
  • Zero-shot LLM-based methods. MapGPT [8] introduces a language-based online map representation to equip GPT-4v [1] with global spatial understanding, enabling adaptive multi-step planning for indoor VLN in a zero-shot setting. STMR[13] enhances spatial reasoning by introducing a Semantic-Topo-Metric Representation that allows LLMs to infer navigation decisions directly from structured scene information. CityNavAgent[45] performs hierarchical semantic planning to predict waypoints in a zero-shot manner, supported by a global memory module that records past trajectories for improved long-horizon navigation.
@@
  • Learning-based methods. LingUNet [27] is an instruction-conditioned visual model adapted from goal-visible aerial navigation to step-wise VLN in AerialVLN[25] Seq2Seq [4] predicts actions using a recurrent policy that encodes visual observations and language instructions into a unified hidden representation. CMA [18] introduces cross-modal attention to better fuse RGB, depth, and textual cues for action prediction. LAG[25] adds a look-ahead guidance module. Grid-VS[48] formulates aerial VLN as a grid-based view selection task and leverages a bird’s-eye-view representation with a cross-modal transformer to model long-term navigation context and instruction alignment. NaVid [44] is a video-based vision-language model that has demonstrated strong performance in indoor VLN. OpenFly [12] adopts a keyframe-aware VLN architecture and uses adaptive token sampling over historical observations.
Method Observation Validation Seen Validation Unseen
[…]
Grid-based VS [48] $\checkmark$ $\checkmark$ 70.3 20.8 33.4 10.2 121.3 7.4 16.1 2.5
CityNavAgent [45] $\checkmark$ $\checkmark$ $\checkmark$ 80.8 13.9 30.2 5.1 60.2 11.7 35.2 5.0
Random Sampling 109.6 0.0 0.0 0.0 149.7 0.0 0.0 0.0
Action Sampling 213.8 0.9 5.7 0.3 237.6 0.2 1.1 0.1
LingUNet [27] $\checkmark$ $\checkmark$ 383.8 0.6 6.9 0.2 368.4 0.4 3.6 0.9
Seq2Seq [4] $\checkmark$ $\checkmark$ 146.0 4.8 19.8 1.6 218.9 2.3 11.7 0.7
CMA [18] $\checkmark$ $\checkmark$ 121.0 3.0 23.2 0.6 172.1 3.2 16.0 1.1
LAG [25] $\checkmark$ $\checkmark$ 90.2 7.2 15.7 2.4 127.9 5.1 10.5 1.4
STMR [13] $\checkmark$ $\checkmark$ 96.3 12.6 31.6 2.7 119.5 10.8 23.0 1.9
MapGPT [8] $\checkmark$ 124.9 2.1 4.7 0.8 107.0 0.0 0.0 0.0
Navid [12] $\checkmark$ 105.1 6.8 15.5 1.1 106.9 6.1 12.3 0.7
OpenFly [44] $\checkmark$ 127.2 8.1 21.8 1.6 113.8 7.6 18.2 1.5
Ours $\checkmark$ 79.6 11.4 37.7 6.3 95.8 8.1 28.9 2.2
@@
IV-B Results on Benchmarks
IV-B1 Quantitative Results on AerialVLN-S Benchmark
To ensure a fair comparison, we categorize methods by their sensor configurations, distinguishing RGB-only agents from those that rely on panoramic inputs, depth sensing, or odometry in Table II. Within the RGB-only group, our method achieves strong performance on all metrics and attains the best overall results, comparable to STMR [13], which required external depth information. The proposed method is also competitive with panoramic baselines, especially in SDTW. This is particularly meaningful in the AerialVLN setting, where many trajectories contain turn-backs and repeated segments, making SDTW a more informative measure of instruction following than SR alone. Panoramic methods typically require multiple cameras or repeated in-place rotations to build 360° observations, which increases time and energy costs for UAVs, whereas our model operates from a single-view RGB stream and thus better matches realistic deployment constraints. MapGPT [8] and Navid [44], which are strong indoor VLN baselines, perform notably worse in outdoor aerial settings, underscoring the additional challenges posed by large-scale, sparsely structured environments. Compared with the most related OpenFly [12] framework, our approach further benefits from long-horizon history modeling and auxiliary tasks, leading to more stable long-range behavior. Overall, these results establish a new state of the art for monocular RGB-only aerial VLN and validate the effectiveness of our unified framework for spatial, temporal, and embodied reasoning. Fig. 6 compares the predicted 3D flight paths with the reference routes. The predicted trajectories preserve the overall geometric structure of the routes, including the major turning segments, long-range direction changes, and the characteristic height variations implied by the instruction. Although small local deviations occasionally appear, the agent often recovers from these offsets and continues to produce geometrically coherent paths over extended horizons. The trajectories also converge toward the intended final region, reflecting stable motion patterns derived solely from egocentric RGB observations in large-scale outdoor environments.
TABLE III: Comparison of different methods on the OpenFly-S test-seen split.
Method Observation OpenFly-S Test Seen
[…]
Random Sampling 242 0.7 0.8 0.0
Seq2Seq [4] $\checkmark$ $\checkmark$ 291 0.1 3.2 0.6
CMA [18] $\checkmark$ $\checkmark$ 144 5.7 30.0 5.1
LAG [25] $\checkmark$ $\checkmark$ 130 6.6 40.8 7.0
Navid [44] $\checkmark$ 142 9.9 24.3 3.9
OpenFly [12] $\checkmark$ 98 33.2 63.5 22.8
Ours $\checkmark$ 38 54.5 72.1 49.9
Figure 7: Comparison between OpenFly and our method across three difficulty levels (Easy, Moderate, and Hard), grouped by trajectory length.
IV-B2 Quantitative Results on OpenFly-S Benchmark
We report quantitative results on the OpenFly-S benchmark in Table III. Our method achieves superior performance compared to baselines across all evaluation metrics, including NE, SR, OSR, and SPL. We further analyze success rates across different difficulty levels defined by trajectory length, following the evaluation protocol adopted in OpenFly [12]. Specifically, episodes with fewer than 30 actions are categorized as Easy, those with 30-60 actions as Moderate, and those with more than 60 actions as Hard. As shown in Fig. 7, the proposed method beats OpenFly[12] across all difficulty levels, with a larger margin on hard episodes. Notably, performance drops for both methods as the trajectory length increases, suggesting that long-horizon aerial navigation remains challenging. We observe that overall success rates on OpenFly-S are higher than those on AerialVLN-S. This behavior is consistent with the characteristics of the OpenFly dataset, which features shorter navigation trajectories and fewer action steps on average. Overall, the experimental results on both AerialVLN-S and OpenFly-S demonstrate that the proposed method maintains strong and consistent performance across multiple aerial VLN datasets constructed under different design principles, supporting its robustness to variations in trajectory scale, instruction length, and data generation strategies.
@@ IV-B2 Quantitative Results on OpenFly-S Benchmark
IV-B3 Cross-dataset Evaluation
Cross-dataset generalization is a critical capability for vision-and-language navigation models. Inspired by the evaluation protocol in ground VLN [44, 10, 32], we assess cross-dataset generalization in aerial VLN by training models on AerialVLN-S and directly evaluating them on AerialVLN without any fine-tuning. Compared to AerialVLN-S, AerialVLN presents a more challenging navigation scenario, with longer path lengths, larger-scale environments, and more complex long-horizon action sequences [25]. As Table IV shown, while all methods exhibit performance degradation under this distribution shift, our method achieves consistently better results across all metrics. These results demonstrate that our method exhibits stronger transferability and generalization capability, even under challenging cross-dataset conditions.We further report results of the proposed model trained without auxiliary tasks. Removing the auxiliary tasks results in inferior performance in the cross-dataset setting, which suggests that these tasks help enhance the agent’s spatial and temporal reasoning over the environment, thereby improving robustness under distributional shifts.
TABLE IV: Cross-dataset Generalization Evaluation on AerialVLN Val-Unseen split. Note that $\dagger$ indicates our method without auxiliary tasks.
Method Observation AerialVLN Validation Unseen
[…]
Seq2Seq [4] $\checkmark$ $\checkmark$ 352.7 0.6 4.2 0.3
CMA [18] $\checkmark$ $\checkmark$ 337.0 1.0 6.7 0.4
Ours $\checkmark$ 311.9 1.8 12.1 0.5
Ours $\checkmark$ 331.5 2.8 13.3 0.8
Figure 9: Qualitative results on auxiliary VQA tasks. Top: examples from the Open3D-VQA dataset [46], where the model answers 3D spatial queries including relative relations and geometric orientation. Bottom: examples from GQA [16], covering object grounding and relational reasoning. These examples illustrate the model’s of spatial understanding and positional reasoning, which is crucial for language-guided UAV navigation.
IV-B4 Qualitative Analysis
@@
TABLE V: Comparison of Model Efficiency.
Method Peak GPU Memory Inference Latency
OpenFly (7B) $\sim$ 14 GB 0.5 s
Ours (8B) $\sim$ 17 GB 0.7 s
Ours (8B)-AWQ $\sim$ 6 GB 0.5 s
IV-C Model Efficiency
We evaluate the inference efficiency of our method to assess its practicality for aerial vision-language navigation. We compare our approach with OpenFly [12], a recent state-of-the-art aerial VLN method, as both methods adopt a VLM-based design paradigm. All measurements are conducted on a single NVIDIA A100 GPU. We report end-to-end inference latency per step  [^1] End-to-end inference latency is measured as the wall-clock time from receiving multimodal observations (RGB images and language instruction) to producing the predicted action. together with peak GPU memory. As shown in Table V, the full-precision 8B model exhibits a per-step inference latency of about 0.7 s and a peak GPU memory usage of approximately 17 GB. We further note that the favorable inference latency of OpenFly [12] mainly benefits from its limited visual input, leading to lower multimodal processing overhead and reduced temporal visual context. To alleviate the memory overhead, we further evaluate an Activation-aware Weight Quantization (AWQ) [22] version of the 8B model. AWQ substantially reduces the memory footprint and lowers the inference latency to around 0.5 s, resulting in a decision rate close to 2 Hz. Further improvements in inference efficiency are expected to benefit from more advanced acceleration techniques and hardware-aware optimizations.
IV-D Ablation Study
TABLE VI: Ablation study on auxiliary tasks: Spatial Perception (SP) and Trajectory Reasoning (TR). Both tasks improve success rate (SR) and trajectory quality (SDTW) in both seen and unseen environments.